# A self-supervised framework for learning whole slide representations

### v1

@@ Abstract
Whole slide imaging is fundamental to biomedical microscopy and computational pathology. However, whole slide images (WSIs) present a complex computer vision challenge due to their gigapixel size, diverse histopathologic features, spatial heterogeneity, and limited/absent data annotations. These challenges highlight that supervised training alone can result in suboptimal whole slide representations. Self-supervised representation learning can achieve high-quality WSI visual feature learning for downstream diagnostic tasks, such as cancer diagnosis or molecular genetic prediction. Here, we present a general self-supervised whole slide learning (S3L) framework for gigapixel-scale self-supervision of WSIs. S3L combines data transformation strategies from transformer-based vision and language modeling into a single unified framework to generate paired views for self-supervision. S3L leverages the inherent regional heterogeneity, histologic feature variability, and information redundancy within WSIs to learn high-quality whole-slide representations. We benchmark S3L visual representations on two diagnostic tasks for two biomedical microscopy modalities. S3L significantly outperforms WSI baselines for cancer diagnosis and genetic mutation prediction. Additionally, S3L achieves good performance using both in-domain and out-of-distribution patch encoders, demonstrating good flexibility and generalizability.
Self-supervised learning, representation learning, histopathology, biomedical microscopy
Figure 1: Self-supervised whole slide learning. Previous work in computational pathology relies on multiple instance learning with weak supervision from slide or patient level labels to learn whole slide representations (Zhu et al., 2017; Ilse et al., 2018; Lu et al., 2021b; Shao et al., 2021; Chen et al., 2022b; Javed et al., 2022; Chen et al., 2023). Here, we present a self-supervised framework for learning whole slide representations, called S3L, by combining data transformations from vision and language modeling to generate high-quality paired views.
@@ 1 Introduction
Whole slide imaging is an integral part of pathology and laboratory medicine. Computational pathology can provide rapid tissue analysis, such as tumor detection, subtyping, and grading, of complex WSIs. Modern computational methods have advanced beyond morphology-based diagnostics to “omics” predictions, prognostication, and treatment response predictions from WSIs alone (Hollon et al., 2023; Coudray et al., 2018; Chen et al., 2022c). Previous results have relied on pathologic annotations for whole slide representation learning; unfortunately, these annotations are weak and expensive to obtain (Campanella et al., 2019), resulting in suboptimal whole slide representations. Few previous studies have focused on learning high-quality whole slide representations using self-supervision (Chen et al., 2022b, 2023; Lazard et al., 2023). Self-supervised whole slide learning can enable better generalization performance to a wide range of downstream pathology tasks with minimal to no annotations required.
In this paper, we present a general and scalable self-supervised framework for whole slide representation learning, S3L. S3L treats gigapixel WSIs as a sequence of patch tokens and applies a domain-informed set of vision-language transformations, including splitting, cropping, and masking, to generate high-quality views for self-supervised training. Over a range of state-of-the-art patch encoders, S3L learns high-quality patch feature aggregation and whole slide representations compared to feature pooling and non-transformer baselines. The main contributions are:
  • We introduce S3L, a general, flexible, and lightweight SSL framework for WSIs at scale and benchmark on two WSI domains and tasks.
  • We design and evaluate the domain-informed vision-language transformations for WSI view generation.
  • We demonstrate that S3L achieves good performance using both in-domain and out-of-distribution patch encoders.
Figure 2: S3L overview. A. The S3L framework consists of a two-stage model architecture: 1) a pre-trained patch encoder $\mathcal{E}$ and 2) a transformer whole slide encoder $f$. WSIs are first divided into small patches, and the patch encoder extracts patch-level features. We then apply whole slide transformations to the patch tokens to create two views of the same WSI. The transformations combine splitting, cropping, and masking, and are informed by the structure and unique properties of WSIs. Each augmented views are encoded by the transformer whole slide encoder, and the feature learning can be self-supervised by any SSL paradigm. B. Example SSL paradigms. In our experiments, we focus on three representative SSL paradigms, including SimCLR, BYOL, and VICReg.
@@ 2.1 Computational pathology
Computational pathology combines whole slide imaging and computer vision methods for the analysis of biomedical tissues. Cancer diagnosis, prognostication, and response-to-treatment are some of the most common computer vision tasks within computational pathology. WSIs pose a unique computer vision challenge due to image sizes ranging up to 150K $\times$ 150K pixels. Data annotations are often limited to whole slide or patient-level labels (Campanella et al., 2019). Moreover, WSIs have a different inherent structure compared to natural images, such as being non-object-centric, containing regional heterogeneity, and visual feature redundancy.
Despite these challenges, significant progress has been made over the last decade due to more accessible slide scanners and modern deep learning methods. Cancer diagnosis (Campanella et al., 2019; Lu et al., 2021a, b; Hollon et al., 2020), molecular classification (Coudray et al., 2018; Hollon et al., 2023), and prognostication (Chen et al., 2021) with WSIs represent how computational pathology can be used to directly improve patient care. Whole slide representation learning for gigapixel WSI search and public WSI-natural language supervision represent future directions in computational pathology (Chen et al., 2022a; Huang et al., 2023).
@@ 2.2 Multiple instance learning
Multiple instance learning (MIL) is a type of supervised learning task where the labels are available only for bags of instances, rather than each individual instance (Dietterich et al., 1997). Learning to classify WSIs has generally been regarded as a MIL task, where each patch is an instance and the WSI is a bag of patches (Ilse et al., 2018). There are two main components in MIL for WSI: the patch encoder and aggregator. Early works directly train the patch encoder in a weakly supervised fashion with WSI labels and generate a whole slide representation with mean pooling (Zhu et al., 2017; Hou et al., 2016). Many works (Campanella et al., 2019; Li et al., 2019; Lu et al., 2021b; Shao et al., 2021) used a ResNet model pretrained on ImageNet as their patch encoder, while there were alternative patch encoder architectures such as MI-FCN (Yao et al., 2019, 2020). (Ilse et al., 2018) first used an attention-based aggregator for WSI classification, and most later works used variations of attention or more powerful transformer-based aggregation models (Li et al., 2019; Yao et al., 2020; Li et al., 2021a; Lu et al., 2021b; Shao et al., 2021; Wang et al., 2022b; Yu et al., 2023; Lin et al., 2023). MCAT (Chen et al., 2021) combined WSI with genomic information for multimodal classification. Additive MIL (Javed et al., 2022) introduced intrinsic interpretability to an attention-based aggregator.
@@ 2.3 Self-supervised representation learning
Self-supervised learning (SSL) is a method of learning representations that does not require labels or annotations. SSL defines pretext learning tasks, such as instance discrimination or reconstruction, as training objectives and is typically evaluated by the performance of the learned representations on downstream tasks (Jaiswal et al., 2020). SSL has shown success in natural images (Chen et al., 2020; Grill et al., 2020; Bardes et al., 2021), natural language processing (Devlin et al., 2018; Mikolov et al., 2013), and multimodal learning (Lu et al., 2023; Huang et al., 2023). SSL has been used for patch encoding in WSIs. DSMIL (Li et al., 2021a) applied SimCLR to learn patch representations. HIPT (Chen et al., 2023) built a general-purpose patch encoder for all types of H&E patches and tasks with self-supervised DINOv2 training. Most recently, PLIP (Huang et al., 2023) leverages histology images and comments posted to Twitter for self-supervised vision-language pretraining.
While there have been attempts at learning whole slide representations in an end-to-end manner with supervision (Hemati et al., 2021), or enhancing whole slide performance with a supervised contrastive loss (Wang et al., 2022a), few works have focused on learning self-supervision beyond patches. HiDisc (Jiang et al., 2023) uses a data hierarchy to learn patch representations that are invariant to slide and patient membership. HIPT (Chen et al., 2022b) uses a self-supervised patch encoder and an encoder for relatively small regions, but weak supervision is used for learning WSI representations.
Giga-SSL (Lazard et al., 2023) and SS-MIL (Tavolara et al., 2022) attempted to learn whole slide representations using self-supervision by masking patch tokens. Both methods have shown limited success compared to their baselines. Giga-SSL also emulates slide augmentation using pre-computed augmented patch representations. However, this strategy does not span all combinations of possible random augmentations, is memory/compute intensive, and is infeasible for large-scale datasets.
@@ 3 Methods
3.1 The S3L framework
WSIs are partitioned into smaller non-overlapping fields-of-view, or patches. Let $\boldsymbol{I}=\left\{I_{i}\right\}_{i=1}^{n}$ be a WSI, where $n$ is the number of patches in a WSI, and each $I_{i}\in\mathbb{R}^{3\times H\times W}$, with coordinate $p_{i}\in\mathbb{Z}^{2}$. $\boldsymbol{p}=\left\{p_{i}\right\}_{i=1}^{n}$ are the coordinates of all patches in the WSI. S3L is a two-stage learning framework to learn WSI representations: 1) a pre-trained patch encoder $\mathcal{E}$; and 2) a transformer-based whole-slide encoder $f$. The pre-trained patch encoder $\mathcal{E}$ can have the architecture of any visual feature extractor. The overall model architecture of S3L is illustrated in Figure 2.
In the S3L framework, the patch encoder $\mathcal{E}$ encodes each $I_{i}$ into a patch token, ${x}_{i}=\mathcal{E}(I_{i})$. We represent all patch tokens in a WSI as $\boldsymbol{x}=\{x_{i}\}_{i=1}^{n}=\mathcal{E}(\boldsymbol{I})$. The whole-slide encoder $f$ serves as an aggregation function that learns a whole-slide representation using patch tokens and their corresponding coordinates: $h=f(\boldsymbol{x},\boldsymbol{p})=f(\mathcal{E}(\boldsymbol{I}),\boldsymbol{p})$. $f$ can be any learned aggregation architecture, such as in ABMIL (Ilse et al., 2018) or a transformer (Vaswani et al., 2017; Shao et al., 2021). Due to the scale of the WSI, it is infeasible to both train the patch encoder $\mathcal{E}$ and whole-slide encoder $f$ jointly in an end-to-end manner. Thus, we freeze the patch encoder $\mathcal{E}$ to allow for large mini-batch training of the whole-slide encoder $f$. The aim of the S3L framework is to learn high-quality whole slide representations with $f$ using self-supervision.
Siamese self-supervised methods share a common strategy: apply two random transformations to a single data example to generate two distinct views, called a positive pair. We designed our S3L framework to be compatible with existing siamese SSL objectives, where a WSI undergoes a transformation into different views: $t(\boldsymbol{I})=[I_{i}\dots I_{k}]$, $t(\boldsymbol{p})=[p_{i}\dots p_{k}]$, $t^{\prime}(\boldsymbol{I})=[I_{j}\dots I_{\ell}]$, and $t^{\prime}(\boldsymbol{p})=[p_{j}\dots p_{\ell}]$, where $t,t^{\prime}\in\mathcal{T}$ are randomly drawn from a set of transformations described in section 3.2. The patches and their corresponding coordinates from these transformed views are processed through $\mathcal{E}$ and $f$ to obtain the whole slide representations:
@@ 3.1 The S3L framework
$\mathcal{L}_{\text{S3L}}\left(h,h^{\prime}\right).$
@@
In our experiments, we selected representative methods from different SSL families as identified by (Balestriero et al., 2023), including SimCLR (Chen et al., 2020) from contrastive learning, BYOL (Grill et al., 2020) from self-distillation, and VICReg (Bardes et al., 2021) from canonical component analysis.
3.2 Whole slide transformations
@@
While previous work focused on pixel space transformations (Lazard et al., 2023), we hypothesized that patch augmentations are alone insufficient to generate high-quality views for whole slide SSL. A patch encoder trained with instance discrimination, for example, should be invariant to pixel-space augmentations. Thus, pixel space augmentations, such as random cropping, should have minimal effect on the representation space given a strong patch encoder. Similar representations across augmentations are especially significant when the patch encoder is trained with a self-supervised loss, as illustrated with SimCLR in Figure 3, because it is explicitly enforced as a pretext task. We opt to bypass patch augmentations altogether, thereby reducing S3L memory and compute burden, and focus on transformations at the WSI level.
@@
Figure 3: Limited effect of pixel-level patch augmentations. We qualitatively evaluate the effect of pixel-level augmentation on the patch representations by visualizing the tSNE plot of SimCLR pretrained patch representations sampled from a single WSI. We observe that strong augmentations at pixel-level have a minimal effect on the patch embeddings. The invariant behavior of the patch encoder is explicitly enforced by the SimCLR pretext task.
Figure 4: S3L transformation strategy. S3L transformation strategy combines splitting, cropping, and masking, and they are informed by the size, region diversity, and information redundancy of WSIs. Splitting partitions patches into mutually exclusive sets decreases mutual information between views; cropping can generate spatially diverse views covering different regions on the WSI; masking reduces redundant visual features and improves training efficiency. The combination of these transformations can create optimal positive pairs for whole slide representation learning.
@@
WSI are gigapixel images divided into a sequence of tokens. Therefore, our whole slide transformations combine strategies from both vision and language modeling, which are selected to address the domain-specific properties of WSI (Figure 4): splitting, cropping, and masking.
@@
  • Splitting to decrease mutual information between views. Reducing mutual information between SSL views, while keeping task-relevant information intact, improves downstream performance (Tian et al., 2020). The splitting transformation randomly partitions patch tokens into disjoint sets, ensuring mutually exclusive set membership for each token. Enforcing mutual exclusion reduces mutual information between SSL views.
@@
  • Cropping to capture regional heterogeneity. WSIs contain regional differences in histologic features. Similar to multi-cropping in visual SSL (Chen et al., 2020; Caron et al., 2020, 2021), cropping generates spatially diverse views of WSI with variable sizes and features.
@@
  • Masking to reduce redundant visual features. Due to their gigapixel size, WSIs can contain large regions of similar histologic features and tissue phenotypes, causing redundancy. Masking is an effective transformation for vision and language modeling (Devlin et al., 2018; He et al., 2022; Chen et al., 2022b). For each view, the masking transformation samples a subset of patch tokens without replacement. Additionally, masking increases training efficiency by reducing sequence lengths.
@@
While splitting, cropping and masking can be used individually, we apply them sequentially to generate versatile and high-quality views for S3L training.
3.3 S3L Implementation

### v2

@@ Abstract
Whole slide imaging is fundamental to biomedical microscopy and computational pathology. Previously, learning representations for gigapixel-sized whole slide images (WSIs) has relied on multiple instance learning with weak labels, which do not annotate the diverse morphologic features and spatial heterogeneity of WSIs. A high-quality self-supervised learning method for WSIs would provide transferable visual representations for downstream computational pathology tasks, without the need for dense annotations. We present Slide Pre-trained Transformers (SPT) for gigapixel-scale self-supervision of WSIs. Treating WSI patches as tokens, SPT combines data transformation strategies from language and vision modeling into a general and unified framework to generate views of WSIs for self-supervised pretraining. SPT leverages the inherent regional heterogeneity, histologic feature variability, and information redundancy within WSIs to learn high-quality whole slide representations. We benchmark SPT visual representations on five diagnostic tasks across three biomedical microscopy datasets. SPT significantly outperforms baselines for histopathologic diagnosis, cancer subtyping, and genetic mutation prediction. Finally, we demonstrate that SPT consistently improves whole slide representations when using off-the-shelf, in-domain, and foundational patch encoders for whole slide multiple instance learning.
Figure 1: Self-supervised whole slide learning. Previous work in computational pathology relies on multiple instance learning with weak supervision from slide or patient-level labels to learn whole slide representations [1, 2, 3, 4, 5, 6, 7]. We present a self-supervised framework for learning whole slide representations, called Slide Pre-trained Transformers (SPT), by combining data transformations from vision and language modeling to generate high-quality paired views.
@@ 1 Introduction
Whole slide imaging is an integral part of tissue diagnosis and laboratory medicine. Computational pathology can provide rapid tissue analysis of complex WSIs, such as cancer detection, subtyping, and grading. Modern computational methods have advanced beyond morphology-based diagnostics to “omics" predictions, prognostication, and treatment response predictions from WSIs alone [8, 9, 10]. Supervised multiple instance learning (MIL) methods, including ABMIL [2], CLAM [3], DSMIL [11], TransMIL [4], can achieve good performance on the above diagnostic tasks. Unfortunately, these methods rely on slide annotations to learn whole slide representations. WSI annotations are weak, sparse, incomplete, and expensive to obtain [12]. Moreover, WSIs are gigapixel-size and contain diverse morphologic and histopathologic features with extensive spatial heterogeneity. Weak slide labels may annotate only a small region within WSIs, demonstrating the limitation of relying on weak supervision to achieve high-quality and transferable whole slide representations.
To compensate for weak whole slide labels, self-supervised learning (SSL) has been increasingly used in computational pathology. The majority of previous SSL work in computational pathology has focused on region or patch learning. HiDisc, PLIP, and UNI are examples of WSI patch encoders that can be used to obtain visual features for downstream whole slide learning [13, 7, 14]. However, few previous studies have investigated learning whole slide representations using SSL [15, 5]. A general and unified framework for whole slide SSL would enable transferable whole slide feature learning and generalize to a wide range of downstream pathology tasks with minimal to no annotations required.
In this paper, we present Slide Pre-trained Transformers (SPT) for self-supervised whole slide representation learning. SPT treats gigapixel WSIs as a sequence of patch tokens and applies a domain-informed set of vision-language transformations, including splitting, cropping, and masking, to generate distinct views for self-supervised training. Over a range of patch encoders, SPT learns high-quality patch feature aggregation and whole slide representations compared to the state-of-the-art baselines. The main contributions are:
  • We introduce SPT, a general, flexible, and unified learning framework for WSIs at scale and benchmark on five computational pathology tasks.
  • SPT outperforms previous state-of-the-art self-supervised and supervised methods for WSI representation learning.
  • SPT offers a consistent performance boost across a wide range of patch encoders.
@@ 2 Related Work
Figure 2: SPT overview. A. The SPT framework consists of a two-stage model architecture: 1) a pre-trained patch encoder $\mathcal{E}$ and 2) a transformer whole slide encoder $f$. WSIs are first divided into small patches, and the patch encoder extracts patch-level features. We then apply whole slide transformations to the patch tokens to create two views of the same WSI. The transformations combine splitting, cropping, and masking, which are informed by the structure and unique properties of WSIs. The transformed views are encoded by the transformer whole slide encoder, and the slide-level feature learning can use any paradigm. B. Example learning paradigms. In our experiments, we focus on three representative self-supervised paradigms, including SimCLR [16], BYOL [17], and VICReg [18], and supervised contrastive learning [19].
@@ 2.1 Computational pathology
Computational pathology combines whole slide imaging and computer vision methods for the analysis of biomedical tissues. Cancer diagnosis, prognostication, and response-to-treatment are some of the most common computer vision tasks within computational pathology. WSIs pose a unique computer vision challenge due to image sizes ranging up to 150K $\times$ 150K pixels. Data annotations are often limited to whole slide or patient-level labels [12]. Moreover, WSIs have a unique data structure compared to natural images, such as being non-object-centric and containing regional heterogeneity and visual feature redundancy. Despite these challenges, significant progress has been made over the last decade due to more accessible slide scanners and modern deep-learning methods. Cancer diagnosis [12, 20, 3, 21], molecular classification [9, 8], and prognostication [22] with WSIs represent how computational pathology can be used to improve diagnostic medicine and patient care. Whole slide representation learning for gigapixel WSI search and public WSI-natural language supervision represent future directions in computational pathology [23, 14].
@@ 2.2 Multiple instance learning
Multiple instance learning (MIL) is a type of supervised learning such that labels are available only for bags of instances, rather than individual instances [24]. Learning to classify WSIs has generally been regarded as a MIL task, where each patch is an instance and the WSI is a bag of patches [2]. There are two main components in MIL for WSI: the patch encoder and aggregator. Early works directly train the patch encoder in a weakly supervised fashion with WSI labels and generate a whole slide representation with mean pooling [1, 25]. Many works have used ResNet or related architectures for patch encoders, often pre-trained on ImageNet [26, 27, 12, 28, 3, 4]. Attention-based multiple instance learning (ABMIL) was first introduced in 2018 for WSI classification [2]. Subsequently, clustering-constrained attention multiple instance learning (CLAM) was used for large-scale weakly supervised computational pathology. Subsequent MIL work has used variations of attention and more powerful transformer-based aggregation models, such as TransMIL [28, 27, 11, 3, 4, 29, 30, 31].
@@ 2.3 Self-supervised representation learning
Self-supervised learning (SSL) is a method of learning representations that does not require labels or annotations. SSL defines pretext learning tasks, such as instance discrimination or reconstruction, as training objectives and is typically evaluated by the performance of the learned representations on downstream tasks [32]. SSL has shown success in natural images [16, 17, 18], natural language processing [33, 34, 35, 36], and multimodal learning [14, 37]. SSL has been used for patch encoding in WSIs. DSMIL [11] applied SimCLR to learn patch representations. Most recently, PLIP [14] leverages histology images and comments posted to Twitter for self-supervised vision-language pretraining. UNI [7] built a general-purpose patch encoder for all types of H&E patches and tasks training with self-supervised DINOv2 [38] on a large-scale institutional dataset.
While there have been attempts at learning whole slide representations in an end-to-end manner with supervision [39], or enhancing them with a supervised contrastive loss [40], few works have focused on self-supervision beyond patches. HiDisc [13] uses a data hierarchy to learn patch representations that are invariant to slide and patient membership. HIPT [5] uses a self-supervised patch encoder and an encoder for local regions, but weak supervision is used for learning WSI representations. Giga-SSL [15] and SS-MIL [41] attempted to learn whole slide representations using self-supervision by masking patch tokens. Giga-SSL also emulates slide augmentations using pre-computed augmented patch representations. However, this strategy is memory intensive, does not span all combinations of possible random augmentations, and is infeasible for large-scale datasets.
@@ 3 Methods
3.1 The SPT framework
WSIs are partitioned into smaller non-overlapping fields-of-view or patches. Let $\boldsymbol{I}=\left\{I_{i}\right\}_{i=1}^{n}$ be a WSI, where $n$ is the number of patches in a WSI, and each $I_{i}\in\mathbb{R}^{3\times H\times W}$, with coordinate $p_{i}\in\mathbb{Z}^{2}$. $\boldsymbol{p}=\left\{p_{i}\right\}_{i=1}^{n}$ are the coordinates of all patches in the WSI. SPT is a two-stage learning framework to learn WSI representations: 1) a pre-trained patch encoder $\mathcal{E}$; and 2) a transformer whole slide encoder $f$. The pre-trained patch encoder $\mathcal{E}$ can have the architecture of any visual feature extractor. The overall model architecture of SPT is illustrated in Figure 2.
In the SPT framework, the patch encoder $\mathcal{E}$ encodes each $I_{i}$ into a patch token, ${x}_{i}=\mathcal{E}(I_{i})$. We represent all patch tokens in a WSI as $\boldsymbol{x}=\{x_{i}\}_{i=1}^{n}=\mathcal{E}(\boldsymbol{I})$. The whole-slide encoder $f$ serves as an aggregation function that learns a whole-slide representation using patch tokens and their corresponding coordinates: $h=f(\boldsymbol{x},\boldsymbol{p})=f(\mathcal{E}(\boldsymbol{I}),\boldsymbol{p})$. $f$ can be any learned aggregation architecture, such as in ABMIL [2] or a transformer [4, 42]. Due to the scale of WSIs, it is infeasible to both train the patch encoder $\mathcal{E}$ and whole-slide encoder $f$ jointly in an end-to-end manner. Thus, we freeze the patch encoder $\mathcal{E}$ to allow for large mini-batch training of the whole-slide encoder $f$. The aim of the SPT framework is to learn high-quality whole slide representations with $f$ using self-supervision.
Self-supervised methods share a common strategy: apply random transformations to a single data example to generate distinct views, called positive pairs. We designed our SPT framework to be compatible with existing SSL objectives, where a WSI undergoes transformations into different views. With a two-view SSL paradigm, $t(\boldsymbol{I})=[I_{i}\dots I_{k}]$, $t(\boldsymbol{p})=[p_{i}\dots p_{k}]$, $t^{\prime}(\boldsymbol{I})=[I_{j}\dots I_{\ell}]$, and $t^{\prime}(\boldsymbol{p})=[p_{j}\dots p_{\ell}]$, where $t,t^{\prime}\in\mathcal{T}$ are randomly drawn from a set of transformations described in section 3.2. The patches and their corresponding coordinates from these transformed views are processed through $\mathcal{E}$ and $f$ to obtain the whole slide representations:
@@ 3.1 The SPT framework
$\mathcal{L}_{\text{SPT}}\left(h,h^{\prime}\right).$
@@
In our experiments, we selected representative methods from different SSL families [43], including SimCLR [16] from contrastive learning, BYOL [17] from self-distillation, and VICReg [18] from canonical component analysis. We denote the self-supervised SPT as ssSPT.
SPT with supervision.
SPT can be adapted to fully supervised training using weak slide- or patient-level labels, by applying a supervised contrastive loss. Positive pairs for supervised contrastive learning are defined by class labels [19]. We denote the supervised variant of SPT as suSPT.
3.2 SPT transformation
@@
While previous work focused on pixel space transformations [15], we hypothesized that these augmentations at patch level are insufficient to generate high-quality views for whole slide SSL. A patch encoder trained with instance discrimination, for example, should be invariant to pixel-space augmentations. Thus, pixel space augmentations, such as color jittering, should have minimal effect on the representation space and resulting similar embeddings, as illustrated with SimCLR in Figure 3, because it is explicitly enforced as the pretext task. We opt to bypass patch augmentations altogether, thereby reducing SPT memory and compute burden, and focusing on transformations at WSI level.
@@
Figure 3: Limited effect of pixel-level patch augmentations. We qualitatively evaluate the effect of pixel-level augmentation on the patch representations by visualizing the tSNE plot of SimCLR pre-trained patch representations sampled from a single WSI. We observe that strong augmentations at the pixel level have a minimal effect on the patch embeddings. The invariant behavior of the patch encoder is explicitly enforced by the SimCLR pretext task.
Transformation strategy.
@@
WSIs are divided into a sequence of patch tokens. The SPT transformation strategy, as shown in Figure 4, is inspired by both vision and language modeling, and was selected to address the domain-specific properties of WSI:
@@
  • Splitting to decrease mutual information between views. Reducing mutual information between SSL views, while keeping task-relevant information intact, improves downstream performance [44, 45]. The splitting transformation randomly partitions patch tokens into disjoint sets, ensuring mutually exclusive set membership for each token. Enforcing mutual exclusion reduces mutual information between SSL views.
@@
  • Cropping to capture regional heterogeneity. WSIs contain regional differences in histologic features. Similar to multi-cropping in visual SSL [16, 46, 47], cropping generates spatially diverse views of WSI with variable sizes and features.
@@
  • Masking to reduce redundant visual features. Due to their gigapixel size, WSIs can contain large regions of similar histologic features and tissue phenotypes, causing redundancy. Masking is an effective transformation for vision and language modeling [5, 34, 48]. For each view, the masking transformation samples a subset of patch tokens without replacement. Additionally, masking increases training efficiency by reducing sequence lengths.
@@
While splitting, cropping and masking can be used individually, we apply them jointly to generate versatile and high-quality views for SPT training.
Figure 4: SPT transformation strategy. SPT combines splitting, cropping, and masking to generate views, and they are motivated by the size, region diversity, and information redundancy of WSIs. Splitting partitions patches into mutually exclusive sets decreases mutual information between views; cropping can generate spatially diverse views covering different regions on the WSI; masking reduces redundant visual features and improves training efficiency. The combination of these transformations can create optimal positive pairs for whole slide representation learning.
3.3 SPT Implementation